# SELF-EXPLAIN: Teaching Large Language Models to Reason Complex Questions by Themselves

### v1

@@ Abstract
Large language models (LLMs) can generate intermediate reasoning steps. To elicit the reliable reasoning, the common practice is to employ few-shot chain-of-thought prompting, where several in-context demonstrations for reasoning are prepended to the question. However, such chain-of-thought examples are expensive to craft, especially for professional domains, and can have high variance depending on human annotators. Therefore, this work investigates whether LLMs can teach themselves to reason without human-crafted demonstrations. We propose SELF-EXPLAIN to generate CoT examples by LLMs inspired by “encoding specificity” in human memory retrieval. We find using self-explanations makes LLMs more confident, more calibrated and less biased when answering complex questions. Moreover, we find prompting with self-explanations can even significantly outperform using human-crafted CoTs on several complex question-answering datasets.
@@ 1 Introduction
Recently, with the appearance of large langauge models (LLMs), some sparks of human-like intelligence are shown. LLMs can learn the underlying patterns from several in-context exemplars, which is known as in-context learning (ICL) (1). In ICL, no weight updates are implemented and LLMs construct answers based on their encoded knowledge following the given in-context demonstrations. On the other hand, Chain-of-Thought (CoT) prompting (31) shows that LLMs can generate rationale while answering the questions. Moreover, not only can LLMs learn to reason from in-context CoT examples, but also generate explanations in zero-shot prompting (13). However, despite the evidence of zero-shot reasoning ability, LLMs generally perform badly without any in-context demonstrations, especially when the tasks are complex. High-quality human-crafted CoT demonstrations, which are usually considered as the gold standard (17), are needed to guide LLMs to reason correctly through ICL.
However, designing and collecting CoT demonstrations for prompting LLMs can be expensive and difficult. On the one hand, for some professional domains such as medicine, physicians or medical students may be needed to produce fine-grained explanations of the answers (20; 32). On the other hand, collecting CoT examples (usually in the form of long sentences) through crowdsourcing can have great uncertainty and noise (29; 7). The style or the form of CoTs may mainly depend on human annotators’ own experience. However, Coda-Forno et al. (3) show that different kinds of expressions of context can actually lead to distinct behaviors of LLMs. Additionally, it is difficult to measure whether those human-crafted CoT examples inform the correct distribution and feature space of the task-specific corpus for effective prompting (18).
Therefore, we are motivated to investigate whether LLMs can teach themselves to reason complex tasks without relying on human-designed CoTs. Our contributions are mainly in three folds: (1) We find LLMs themselves can produce explanations of complex questions given the answer, which we refer to SELF-EXPLAIN. (2) Prompting with self-explanation  [^1] We use CoT and explanation interchangeably to refer to sentences showing reasoning in the prompt. through ICL also makes LLMs more calibrated and less biased. For example, our experiments show self-explanation makes LLM 1.2% less biased in multiple-choice questions. (3) Prompting with self-explanation can outperform prompting with human-crafted CoT on challenging reasoning tasks even for clinical QA which requires professional medical knowledge. Our findings may challenge the common wisdom that human-crafted CoTs are the gold standard for prompting LLMs.
2 Related Work
In-Context Learning.
In-Contex Learning (ICL) is the ability of language models to induce answers from given demonstrations without weights updating in supervised tuning. In-context exemplars are the key to ICL which have dominating influence on the generation. Quite a few works have been proposed to optimize the selection of exemplars (16; 24; 4). On the other hand, in the cases of no access to task labels, Lyu et al. (18) proposed zero-shot ICL that employs pseudo exemplars with random labels for classification tasks. Such zero-shot method can be supported by the finding of Min et al. (19) on classification tasks that the label space carries more weight than correctness of labels in exemplars for ICL.
Chain-of-Thought Prompting.
Prompting with reasoning triggers LLMs to generate intermediate steps of thinking, known as Chain-of-Thought (CoT) (31). Some works (22; 12; 21; 35) on multi-step reasoning prompt LLM to decompose a complex question into sub-tasks iteratively, showcasing the LLMs’ capability of compositional reasoning under some human guidance. Additionally, Kojima et al. (13) propose zero-shot prompting to elicit LLMs’ chain-of-thought reasoning without any human-crafted exemplars. This prompting method is leveraged to generate CoT exemplars for ICL (28; 33; 2). Wan et al. (28); Zhang et al. (33) mainly focus on selecting CoTs generated through Kojima et al. (13). Chen et al. (2) further incorporate pseudo task generation apart from pseudo label/ CoTs. Different from those works, we explore a new zero-shot prompting fashion inspired by encoding specificity in human’s memory retrieval. Additionally, we further rigorously investigate the difference between human-crafted CoTs and self-generated CoTs and the according influence on model confidence.
@@ 3 Revisiting ICL
We first detail some annotations and give a formal setup of ICL. We denote the model parameters as $\theta$, CoT as $\pi$ and assume a labeled dataset $\mathcal{D}$ with distribution $p^{\ast}$. Given a test query $\mathbf{x}$, the model will predict $\tilde{\mathbf{y}}$ by conditioning on the query and in-context exemplars. We can then have,
@@ 3.1 Encoding Specificity Hypothesis
A key question for CoT prompting is how to compose in-context CoTs for some specific dataset $\mathcal{D}_{(x,y)}$? CoTs can be rephrased differently while delivering the same logic. To understand this question, we take a memory view of ICL by conceptualizing LLMs as memory networks (6; 11; 34; 10; 23; 14). The feed-forwarding through hidden layers of LLM is to retrieve and generalize learnt information in memory to construct the output $\mathbf{y}$ to complete the query $\mathbf{x}$ under the guidance of context $\mathbf{C}$ (i.e., in-context exemplars). The pretraining stage can be viewed as encoding information into the weights, i.e., memories of LLM.
@@
From a memory view, CoTs should be composed in a way to match the context in the training corpus seen during encoding similar information to the test data. Such requirement on the match of context between testing and training is known as encoding specificity (27) in humans’ memory retrieval. To see this hypothesis, a simple thought experiment can be considered: supposing that a specific datapoint $(x,y)$ has been seen during language modeling in pretraining and $C$ is the corresponding context prepending $(x,y)$, (i.e., a consecutive string $(C,x,y)$ is seen by LLM during training), at test time, prompting the LLM with $(C,x)$ will trigger the LLM to generate $y$ with great probability. All in all, the encoding specificity hypothesis implies in-context exemplars should match the distribution of context seen during pretraining, i.e., the training corpus encoded by LLMs containing similar information to $(x,y)$. In that way, LLMs are expected to be motivated to search from related encoded knowledge to form the output generation.
Figure 1: The overall framework of SELF-EXPLAIN. We first generate self-explanations on some sampled training data. At test time, those self-explanations will be used as in-context exemplars for prompting.
4 SELF-EXPLAIN
Motivated by the encoding specificity, we would like our CoTs to match the training corpus similar to the test data distribution. In order to achieve this, we propose SELF-EXPLAIN where we directly prompt the LLM to elicit the explanation. The overall pipeline of our approach is shown in Fig. 1. The LLM will leverage the encoded knowledge it deems related to the provided question so as to finish the according generation. Those self-explanations are then used as in-context CoT exemplars for ICL. Formally, we consider a setting where we have some data $(x,y)$ sampled from the distribution $p^{\ast}$ and $p^{\text{test}}\approx p^{\ast}$. We show $y$ does not have to be ground-truth in Sec. 5.2. We assume no human-crafted CoTs are available and define the best CoT as $\pi^{\ast}:=\text{argmax}_{\pi}P(y|x,\pi,\theta)$. Self-explanation is represented as,
@@ 4 SELF-EXPLAIN
where $\gamma$ is an instruction. We hope LLMs to generate reasoning path based on given $(x,y)$ by recalling from its according encoded knowledge so as to satisfy encoding specificity. Moreover, we find $P(y|x,\pi^{\text{self}},\theta)>>P(y|x,\pi^{\text{human}},\theta)$ and $P(y|x,\pi^{\text{self}},\theta)\approx 1$. Further discussion with supporting experimental results is in Sec. 6. We may arguably state that $\pi^{\text{self}}$ is a very close estimation to $\pi^{\ast}$ and have the following approximation,
$\text{argmax}_{\pi}P(\pi|x,y,\gamma,\theta)\approx\text{argmax}_{\pi}P(y|x,\pi,\theta).$ (3)
This shows the self-explanation to a specific $x$, is actually more effective and acceptable for LLM to generate the according $y$, which may evidence the encoding specificity hypothesis. Because self-explanation may be generalized from some related training data in a similar distribution to $x$.
@@
Generated self-explanation will then be used as in-context demonstrations together with corresponding question-answer pairs. However, the underlying logic of $\pi^{\text{self}}$ may be very specific to its corresponding $(x,y)$ and thus lacks generalizability to other different cases. Then, the output explanation $\hat{\pi}_{\text{te}}$ at test time may fail to apply to the input cases, leading to wrong answers. To mitigate this issue, we design $\gamma$ so as to prompt the model to generate solutions employing distinct logics. Formally, we have,
$(\pi^{\text{self}}_{1},...,\pi^{\text{self}}_{n})=\text{argmax}_{\pi}P(\pi|x,y,\gamma,\theta),$ (4)
where $n\in(1,N)$ and NN is the number of different explanations to generate. Then $\pi^{\text{self}}_{i}$ for some $(x_{i},y_{i})$ as an in-context exemplar will be randomly sampled from the according $\{\pi^{\text{self}}_{n}|n\in(1,N)\}$ of $(x_{i},y_{i})$ for unseen test cases.
@@ 5.1 Experimental Setup
We are focused on knowledge-intensive question-answering tasks that require logical thinking on information and associating encoded knowledge but mere facts retrieval. Such knowledge-intensive QA is common and important for the applications of LMs (9; 26). We use challenging MedMCQA (20) and MedQA (8) with standard splits that involve multiple-choice questions to diagnose clinical cases, which are used for physician qualification exams. We also evaluate SELF-EXPLAIN on StrategyQA (5) for reasoning in general domains. In terms of implementations, we use GPT-3.5-turbo as our base LLM. Except for Auto-CoT (33), we randomly sample five exemplars from training data and keep exemplars the same for different baselines using ICL. For MedQA which does not provide human-crafted CoTs, in all cases, we use all the five exemplars provided and annotated in Singhal et al. (25).
@@
No CoT 51.7 55.1 46.8
Zero-shot CoT (13) 51.1 60.6 45.6
Auto-CoT (33) 52.5 61.9 52.7
Human CoT 53.1 61.7 56.1
Self-Exp 56.6 63.7 59.7
Table 1: Test accuracy of different prompting methods on three datasets.
5.2 Results
As is shown in Table 1, using CoT prompting gives better performance than using input-output pairs solely. This demonstrates the importance of CoT in complex reasoning tasks to LLMs. More importantly, prompting with our proposed self-explanation can impressively outperform using CoTs crafted by human professionals by 2% in challenging Medical datasets. Considering the difficulty and expense of crafting CoTs by humans, employing SELF-EXPLAIN can thus be very useful in some domains. Our approach also outperforms Auto-CoT (33) and vanilla zero-shot CoT (13). Additionally, our results may challenge the common belief that human-crafted CoTs are gold standards which are expected to give better performance than using machine-generated CoTs.
MedMCQA MedQA StrategyQA
Right 56.6 63.7 59.7
Wrong 56.0 63.5 59.1
Does the correctness of self-explanation matter?
A natural question raised in SELF-EXPLAIN is what if the self-generated CoTs are wrong since the generation process is not supervised by humans. We look into this question by guiding LLMs to generate misleading explanation. More specifically, we provide the LLM with wrong answers to generate corresponding self-explanations. Those self-explanation for wrong answers are then used for prompting as in-context exemplars. The results are shown in Table 2. We find that the performance of prompting with self-explanation is insensitive to its correctness. We speculate that what carries more weight is the underlying distribution of self-explanations. This may evidence encoding specificity hypothesis that the matched distribution of exemplars at test with that of the context seen during encoding relevant information at pre-training is more important for effective prompting through ICL. This result suggests that a labeled dataset is not necessary to elicit CoTs in SELF-EXPLAIN.
Figure 2: Similarity between human-crafted CoTs and self-explanations in terms of ROUGE-L score, terminology covered (terms are extracted by SciSpacy) and length on MedCQA.

### v3

# Large Language Models are In-context Teachers for Knowledge Reasoning

@@ Abstract
In this work, we study in-context teaching (ICT), where a teacher provides in-context example rationales to teach a student to reason over unseen cases. Human teachers are usually required to craft in-context demonstrations, which are costly and have high variance. We ask whether a large language model (LLM) can serve as a more effective in-context teacher for itself or other LLMs, compared to humans. Inspired by the Encoding Specificity Hypothesis from human episodic memory, we hypothesize that in-context exemplars crafted by the teacher should match the training data of the student. This hypothesis motivates us to propose Self-Explain where an LLM’s self-elicited explanations are used as in-context demonstrations for prompting it as they are generalized from the model’s training examples. Self-Explain is shown to significantly outperform using human-crafted exemplars and other baselines.
Furthermore, we reveal that for ICT, rationales from different teacher LLMs or human experts that more resemble the student LLM’s self-explanations are better in-context demonstrations. This supports our encoding specificity hypothesis. We then propose Teach-Back that aligns a teacher LLM with the student to enhance the ICT performance. For example, Teach-Back enables a 7B model to teach the much larger GPT-3.5 in context, surpassing human teachers by around 5% in test accuracy on medical question answering.
Large Language Models are In-context Teachers
for Knowledge Reasoning
Jiachen Zhao1, Zonghai Yao2, Zhichao Yang2, Hong Yu2, 1Northeastern University, 2University of Massachusetts Amherst zhao.jiach@northeastern.edu
Figure 1: Overview of our approaches. Existing few-shot CoT prompting methods rely on human experts to craft rationales as in-context demonstrations. We propose Encoding Specificity Hypothesis to make large language models better in-context teachers than humans. We accordingly design Self-Explain for an LLM itself to be the teacher and Teach-Back to improve the LLM’s capability of teaching another model in context.
@@ 1 Introduction
Knowledge reasoning, different from numerical reasoning, requires large language models (LLMs) to deduce the association between questions and answers that do not usually explicitly appear in the training corpus, although LLMs may have memorized all the facts involved in the question. Such a compositionality gap (Press et al., 2023) between testing and pretraining makes knowledge reasoning difficult and beyond mere fact retrieval. However, LLMs have demonstrated impressive knowledge reasoning performance on diverse tasks (Wei et al., 2022) with few-shot prompting. Exemplars of reasoning are provided in the prompt as context to teach LLMs to reason through in-context learning (ICL) (Brown et al., 2020) at inference. LLMs will generate intermediate reasoning steps (known as Chain-of-Thought (CoT) [^1] We use “CoT” and “rationale” interchangeably to refer to reasoning paths. (Wei et al., 2022)) for deducing the test cases.
Standard few-shot CoT prompting requires humans to first craft high-quality demonstrations of reasoning for LLMs, as depicted in the upper part of Figure 1. However, this may bring some issues. On the one hand, in professional domains such as medicine, experts like physicians are needed to produce fine-grained rationales with correct jargon, which is time-consuming and expensive (Pal et al., 2022; Yang et al., 2023). On the other hand, different from labels, rationales can be phrased in varied ways, while all being correct (Yao et al., 2023). Collecting reasoning examples through crowd-sourcing can thus have great uncertainty (Gebreegziabher et al., 2023). The constructed rationales heavily depend on human annotators’ own experience and thus, may be very subjective (Lee et al., 2022).
More fundamentally, there is a limited understanding of the principles behind constructing effective rationale exemplars for in-context learning. Currently, the majority of works depend on human-crafted demonstrations (usually by professionals) that are based on some heuristic rules (Fu et al., 2023b; Zhou et al., 2022; Khot et al., 2023). However, it is unclear whether those sophisticated rationales crafted by humans are equally the most sensible to LLMs. Demonstrations of rationales from humans may not always be helpful (Yao et al., 2023), although they are often assumed to be gold standards (Muller et al., 2021).
Therefore, we are motivated to ask, can an LLM teach itself or other models through in-context learning for knowledge reasoning, preferably better than humans? We consider a generic framework of in-context teaching (ICT), where a teacher (e.g., human or LLM) constructs example rationales that are then used as in-context demonstrations to prompt a student LLM.
In terms of how to construct exemplars for effective ICT, we consider the Encoding Specificity Hypothesis (Tulving, 1972), which is initially proposed for retrieval from human episodic memory. The hypothesis postulates that the context during recalling information from episodic memory should match the context during encoding. As inspired by the convergence between memory and attention module in language models (Ramsauer et al., 2020; Bricken and Pehlevan, 2021; Zhao, 2023), for few-shot prompting, we similarly hypothesize that in-context exemplars at test time should match the encoded rationales from the student’s training corpora related to the test domain. For example, when the student reasons over medical questions at inference, in-context rationales are expected to be phrased similarly to examples in the medical corpus learned by the student model during training.
The encoding specificity hypothesis can be easily satisfied when the teacher model is the student model itself. We directly prompt the student model to explain the given answer to a question sampled from the same dataset as the test data as inspired by learning theory in cognitive science (Chi et al., 1989). Those elicited self-explanations can represent the model’s encoded knowledge for the test task and are then used as in-context demonstrations at inference. We refer to this approach as Self-Explain. On the other hand, when the teacher model is different from the student model (e.g., using a weak and small model to teach a much larger model (Burns et al., 2024)), we first let the teacher model learn from the student’s self-explanations before eliciting the teacher’s explanations (see the lower part of Figure 1). We refer to this method as Teach-Back, which is how healthcare providers (i.e., teachers) reduce the communication gap with patients (i.e., student) for effective health education (Talevski et al., 2020).
Our experimental results provide sources of evidence for our encoding specificity hypothesis. We find that the student model itself tends to be the best in-context teacher for it, surpassing human teachers or other LLM teachers (w/o Teach-Back). Our experiments across models of different sizes and reasoning abilities suggest that for ICT, larger and stronger models are not necessarily better in-context teachers, though they may produce more reliable rationales. We show that in-context reasoning examples that more resemble the student’s self-explanations can lead to better student performance. This also supports our encoding specificity hypothesis. Furthermore, applying Teach-Back can significantly improve the ICT capability of a teacher model for the student model and even outperform Self-Explain depending on the teacher model. For example, Teach-Back enables a small deployable 7B model to teach the much larger GPT-3.5 in context, surpassing human teachers by around 5% in test accuracy on medical question answering.
In summary, our contributions are mainly in three folds:
  • We investigate in-context teaching for knowledge reasoning, where a teacher provides in-context demonstrations to teach a student model to reason. We propose the Encoding Specificity Hypothesis as the guideline for composing in-context exemplars and provide sufficient evidence for our hypothesis.
  • We propose a new way of eliciting rationales from an LLM by prompting it to explain question-answer pairs. We then propose Self-Explain prompting to use an LLM’s self-explanations as its in-context exemplars, which outperforms using human-crafted CoTs.
  • Our experiments suggest that in-context exemplars of rationales from LLM teachers or human teachers that more resemble the student’s self-explanations may produce better reasoning performance. We then propose Teach-Back, demonstrating that the in-context teaching ability of LLMs can be improved by aligning the teacher with the student’s self-explanations.
Figure 2: The overall framework and prompting format of our approach. The teacher LLM is prompted to generate explanations on sampled training data. Those teacher’s explanations are used as in-context demonstrations for the student model at test time. The student model and the teacher model can be the same.
@@ 2 Revisiting ICL
We first detail some annotations and give a formal setup of in-context learning. We denote the model parameters as $\theta$, rationale as $\pi$ and assume a labeled dataset $\mathcal{D}$ with distribution $p^{\ast}$. Given a test query $\mathbf{x}$, the model will predict $\tilde{\mathbf{y}}$ by conditioning on the query and in-context exemplars. We can then have,
@@ 2.1 Encoding Specificity Hypothesis
A key question for ICL is how to compose in-context rationales for some task dataset $\mathcal{D}_{(x,y)}$? Rationales can be rephrased differently while delivering the same logic. To understand this question, we take a memory view of ICL by conceptualizing LLMs as memory networks Hopfield (1982); Kanerva (1988); Kaiser and Bengio (2016); Ramsauer et al. (2020); Krotov and Hopfield (2016). The feed-forwarding through hidden layers of LLM is to retrieve and generalize learned information in memory to construct the output $\mathbf{y}$ to complete the query $\mathbf{x}$ under the guidance of context $\mathbf{C}$ (i.e., in-context exemplars). The pretraining stage can be viewed as encoding information into the weights, i.e., memories of LLMs.
@@
From a memory perspective, we draw inspiration from the encoding specificity hypothesis Tulving and Thomson (1973), which suggests that successful retrieval of information depends on the match between the context during encoding and the context at retrieval. To see this hypothesis, a simplified thought experiment can be considered: supposing that a specific datapoint $(x,y)$ has been seen during language modeling in pretraining and $C$ is the corresponding context prepending $(x,y)$, (i.e., a consecutive string $(C,x,y)$ is seen by LLM during training), at test time, to let the model generate $y$ with great probability, we can prompt it with $(C,x)$.
More generally, the encoding specificity hypothesis suggests that in-context exemplars of reasoning at test time should match the distribution of reasoning examples seen during training, especially the training corpus containing information similar to task data $\mathcal{D}_{(x,y)}$. LLMs may have seen many sentences involving reasoning during pre-training and further instruction fine-tuning. It may be easier for LLMs to generalize from in-context exemplars similar to those rationales from training data (e.g., having similar reasoning logic or using similar expressions/ jargon) to answer new questions at inference.
3 Methodology
The general framework of our proposed methods and our prompting format is shown in Figure 2. We first introduce Self-Explain, where the student and teacher are the same (Section 3.1), as a straightforward implementation of the encoding specificity hypothesis. We then extend this approach to employing a different teacher model (Section 3.2).
3.1 Self-Explain
Motivated by the encoding specificity, we would like our in-context exemplars of reasoning to match the LLMs’ training corpus containing information similar to the task data. To achieve this, we directly prompt an LLM to elicit its explanation for some question-answer pairs of task data. Such self-explaining is actually how humans integrate new information with their existing knowledge (Chi et al., 1989). Similarly, the LLM is expected to utilize its existing encoded knowledge relevant to the unseen question provided, in order to generate its explanations. These self-explanations are then used as in-context exemplars of reasoning to prompt the model itself.
Eliciting LLMs’ Self-explanations.
Formally, we assume access to labeled training data where we have some data $(x,y)$ sampled from the distribution $p^{\text{train}}$ and assume $p^{\text{test}}\approx p^{\text{train}}$. We consider a realistic setting where human-crafted CoTs are not available. We define an oracle CoT as
$\pi^{\ast}:=\text{argmax}_{\pi}P(y|x,\pi,\theta).$ (2)
Self-explanation is then obtained as,
@@ 3.1 Self-Explain
where $\gamma$ is an instruction. We hope LLMs generate rationales based on a given $(x,y)$ by recalling relevant encoded knowledge so as to satisfy the encoding specificity. We find $P(y|x,\pi^{\text{self}},\theta)>>P(y|x,\pi^{\text{human}},\theta)$ (see Appendix C). We may arguably state that $\pi^{\text{self}}$ is a more reasonable estimation to $\pi^{\ast}$ than $\pi^{\text{human}}$.
Filtering Self-explanations.
We filter out the elicited self-explanations based on the explanation faithfulness (Jacovi and Goldberg, 2020). Explanations that fail to guide the model to produce the given answer, i.e., $y\neq\text{argmax}_{\tilde{y}}P(\tilde{y}|x,\pi^{\text{self}},\theta)$, are screened. We empirically verify the self-explanation ability of different LMs and show that those models succeed in justifying the given $(x,y)$ most of the time (see Section 5.1).
ICL with Self-explanations.
The self-explanations $\pi^{\text{self}}$ elicited by the model are then used as in-context exemplars for prompting it following Equation 1. This can be viewed as the model teaching itself via ICL to do reasoning. Additionally, the ICL performance is very close when using respective $\pi^{\text{self}}$ elicited with either wrong or ground-truth $y$ for the input question $x$ in Equation 3 (see Section 4.2).
@@
The underlying logic of $\pi^{\text{self}}$ might be very specific to its corresponding $(x,y)$ and thus lacks generalizability to other different cases from test data. Then, the output explanation $\hat{\pi}_{\text{te}}$ at test time may fail to apply to the input cases, leading to wrong answers. To mitigate this issue, we also design a new instruction $\gamma^{\prime}$ so as to prompt the model to generate solutions employing distinct logics. Formally, we have,
$(\pi^{\text{self}}_{1},...,\pi^{\text{self}}_{n})=\text{argmax}_{\pi}P(\pi|x,y,\gamma^{\prime},\theta),$ (4)
where $n\in(1,N)$ and $N$ is the number of different explanations to generate. For example, if $N=5$, $\gamma^{\prime}$ will be “Explain how to reach this answer in five different ways”. Then at test time, $\pi^{\text{self}}_{i}$ for an in-context exemplar $(x_{i},y_{i})$ will be randomly sampled from the according $\{\pi^{\text{self}}_{n}|n\in(1,N)\}$ of $(x_{i},y_{i})$.
3.2 Teach-Back
Instead of ICL with self-explanations where the student model teaches itself to reason, these explanations can be provided as in-context exemplars by a different model (parameterized by $\theta_{\text{teacher}}$), i.e.,
$\pi^{\text{self}}=\text{argmax}_{\pi}P(\pi|x,y,\gamma,\theta_{\text{teacher}}).$ (5)
However, explanations of one model may not be the most helpful reasoning demonstrations for another model, especially when the teacher’s explanations are very distinct from the student’s self-explanations (see results in Section 4.3). Based on our encoding specificity hypothesis, we propose to let the teacher model learn from the student’s self-explanations (through supervised fine-tuning) before eliciting the teacher’s explanations. This method is called Teach-Back, which is similar to how healthcare providers reduce the communication gap with patients for effective health education (Talevski et al., 2020). Doctors will rephrase and clarify their explanations based on patients’ explanations for better communication. In Section 4.4, we empirically show the effectiveness of Teach-Back in improving teaching efficacy and enhancing student’s performance.
[…]
No CoT 51.7 52.1 46.8
Zero-shot CoT (Kojima et al., 2022) 51.1 54.4 45.6
Auto-CoT (Zhang et al., 2023) 52.5 55.2 52.7
Human CoT 53.1 55.6 56.1
[…]
w/ Multi-Exp 56.6 59.6 59.7
Table 1: Test accuracy of different prompting methods on three datasets for knowledge reasoning.
@@ 4.1 Experimental Setup
Datasets.
We focus on knowledge-intensive question-answering tasks that require logical reasoning and associating encoded knowledge, rather than just retrieving facts. Such knowledge-intensive QA is common and important for the applications of LLMs Jin et al. (2021b); Tran et al. (2023). We evaluate our method in both general domains and expert domains. We employ widely-used StrategyQA Geva et al. (2021) for commonsense reasoning. For expert domains, we use challenging MedMCQA Pal et al. (2022) and MedQA Jin et al. (2021a) with standard splits. These datasets consist of multiple-choice questions to diagnose clinical cases, which are used for physician qualification exams.
Models.
We use a variety of language models. We employ the chat version of 7B model and 13B model of Llama2 (Touvron et al., 2023), the 7B model of Mistral (Jiang et al., 2023), the Phi3-128k-mini that has 3.8B parameters Abdin et al. (2024) and the frozen version (0613) of GPT-3.5  [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo.
Prompting.
For the instruction used for eliciting models’ self-explanations, an ablation study is conducted in Appendix A. For few-shot prompting at test time, we use five in-context exemplars sampled from the training data.
Baselines.
Apart from comparing our approach with standard few-shot prompting with human CoTs, we include three more baselines. (1) “No CoT”: We remove rationales and use input-output pairs only for in-context exemplars; (2) ”Zero-shot CoT” (Kojima et al., 2022): This method does not require human-crafted demonstrations as it is not few-shot prompting. It directly elicits reasoning from LLMs for the test question by using the prompt “Let’s think step by step”. (3) “Auto-CoT” (Zhang et al., 2023): This work uses the same method as Kojima et al. (2022) to elicit rationales from LLMs. But it further proposes a way of exemplar selection to choose elicited rationales as in-context exemplars. For fair comparison, in each trial, we use the same question-answer pairs for few-shot demonstrations for all baselines.
MedMCQA MedQA StrategyQA
Right 56.6 59.6 59.7
Wrong 56.0 59.4 59.1
Figure 3: The test performance with respect to the number of self-explanations to generate for each exemplar.
@@
No CoT 28.4 31.1 31.8 49.5 41.1
Human 27.3 31.4 38.2 53.3 55.6
[…]
Llama2-13B 30.2 35.5 41.1 55.3 56.9
Mistral-7B 25.1 34.7 44.2 54.4 53.5
Phi3-mini 18.7 35.1 40.7 57.1 57.1
GPT-3.5 18.1 34.4 43.1 57.7 57.5
Table 3: Results of teaching student LLMs with teachers’ self-explanations through in-context learning. The best test accuracy is highlighted in bold.
Figure 4: (a) ROUGE scores between self-explanations of teacher and student. For “Human” teachers, human-crafted CoTs are used for computation. (b) Strong linear correlation is observed between ROUGE scores of self-explanations of teacher and student and the student’s test accuracy.
4.2 Few-shot Prompting with Self-explanations
In this section, we evaluate the test performance of Self-Explain, i.e., when the model’s self-explanations are used as in-context exemplars of reasoning for few-shot prompting. We conduct experiments with GPT-3.5-turbo on reasoning tasks in both general domains and expert domains.